A car approaches a stop sign going 5 m/s before it comes to a stop. If it does this in 4 seconds, what was its acceleration?


1
Expert's answer
2022-03-10T18:00:47-0500

The acceleration of the car

"a=\frac{\Delta v}{\Delta t}"

"a=\rm\frac{0\: m/s-5\: m/s}{4\: s}=-1.25\: m/s^2"